Kevin has $6.45 in coins in his cash box. The numbers of quarters is one less
than twice the number of dimes. The number of nickels is one less than twice the number of quarters. The value of the pennies is the same as the value of the nickels. How many of each Type of coin does he have?

Answer by stanbon(75887)

Kevin has $6.45 in coins in his cash box.
The numbers of quarters is one less than twice the number of dimes.
The number of nickels is one less than twice the number of quarters.
The value of the pennies is the same as the value of the nickels.
How many of each Type of coin does he have?
----
Equations:
p + 5n + 10d + 25q = 645 cents
q = 2d-1
n = 2q-1
p = 5n
-----
Rearrange:
p + 5n + 10d + 25q = 645
0 + 0 + 2d - q = 1
0 + n + 0 - 2q = -1
p - 5n + 0 + 0 = 0
----------------------------
Use any method you know to solve to get:
p = 125
n = 25
d = 7
q = 13
==============
Cheers,
Stan H.